OpenStudy (anonymous):

I need help on simplifying a rational expression p^3+p^2q-2pq^2 divided by pq^2+p^2q-2p^3

OpenStudy (anonymous):

\[\frac{p^3+p^2q-2pq^2}{pq^2+p^2q-2p^3}=\frac{p(p^2+pq-2q^2)}{p(q^2+pq-2p^2)}=\frac{p^2+pq-2q^2}{q^2+pq-2p^2}\] Try factoring

OpenStudy (anonymous):

thats what I need help on I don't know how to factor it

OpenStudy (anonymous):

\[\begin{align*}(p-q)(p+2q)&=p(p+2q)-q(p+2q)\\&=p^2+2pq-pq-2q^2\\&=p^2+pq-2q^2\end{align*}\] The denominator can be factored almost exactly the same way: \[(q-p)(q+2p)=q^2+pq-2p^2\] So \[\frac{p^2+pq-2q^2}{q^2+pq-2p^2}=\frac{(p-q)(p+2q)}{(q-p)(q+2p)}=-\frac{(p-q)(p+2q)}{(p-q)(q+2p)}=-\frac{p+2q}{q+2p}\]

OpenStudy (anonymous):

alright thnaks for the help it was really helpful

OpenStudy (anonymous):

yw